OpenStudy (anonymous):

120.05 cubic centi = ___ mL

OpenStudy (tkhunny):

Have you considered converting it?

OpenStudy (youngjuvie):

are u tryning to covert it to a fraction Convert to a fraction by placing the decimal number over a power of 10

OpenStudy (youngjuvie):

\[\frac{ 2401 }{ 20 }\]

OpenStudy (youngjuvie):

@Glisten

OpenStudy (youngjuvie):

12005%

OpenStudy (tkhunny):

Why are we converting to percentages? \(mL = cm^{3}\) \(120.05\;cm^{3}\cdot\dfrac{1\;mL}{1\;cm^{3}}\) Reconcile the units.

OpenStudy (youngjuvie):

Substitute the value of the variable into the expression and simplify

OpenStudy (anonymous):

i need to cconvert it to cmL

OpenStudy (tkhunny):

No, you don't. That conversion is 1:1. They are the same.

OpenStudy (youngjuvie):

304.9270cm

OpenStudy (youngjuvie):

you have to round up so 307.3400cm

OpenStudy (youngjuvie):

121=3073400 cm

OpenStudy (tkhunny):

What? This is mystery physics. \(120.05\;cm^{3} = 120.05\;mL\) -- Done. Why are the numbers changing? Why are we dividing or rounding? Leave it alone!